A photon has a momentum of magnitude 8.24*10^-28kgm/sec
What is the energy of this photon ?Give your answer in joules and in electro volts.

Grade:

1 Answers

Madalitso Chitsakamile
5 years ago
P = h/lambdaLambda = (6.63*10^-34J/S)/(8.24*10^-28kgm/s)Lambda = 8.046*10^-7mE=hc/lambdaE= 2.47*10^-19J answer 2.47*10^-19/1.6*10^-19 = 1.5eV answer